OpenStudy (anonymous):

(HELP PLEASE) Sally invests $4000 in a savings account which compounds interest monthly. After 10 years, she has a balance of $4,451.52. What is the annual interest rate of the account, rounded to the nearest 0.01%? (please help none of my notes covered this type of problem.)

OpenStudy (jwang55):

P=PO(1+r/k)^kt p=future amount po=initial amount r=rate of change(unknow) k=# of compounding period t=time so 4451.52=4000(1+(x/12))^(12*10) then find out x

OpenStudy (jwang55):

Hope if this makes sense

OpenStudy (anonymous):

Am I suppose to get a negative number?

OpenStudy (jwang55):

no, but it should be super small

OpenStudy (jwang55):

i got 12.84%

OpenStudy (jwang55):

for the annual interest rate

OpenStudy (anonymous):

I was way off I got a negative around -23.78, I must have done something wrong.

OpenStudy (jwang55):